# Understanding the Cluster LP for Correlation Clustering

### v2

@@ 6.3 Wrapping up: Deferred proofs of Lemma 37, 38, 39
Given cluster LP solution $z_{S}$ and $y$, for each node $u$, the covariance matrix $COV_{u}\in\mathbb{R}^{V\times V}$ where
$COV_{u}(v,w)=y_{uvw}-y_{uv}y_{uw}$
is positive semi-definite(PSD). Actually, we can even show a stronger version of the covariance constraint.
@@
$\geq\sum_{S\supseteq T,S^{\prime}\supseteq T}z_{S}z_{S^{\prime}}a^{2}(S)-\sum_{S\supseteq T}\sum_{S^{\prime}\supseteq T}z_{S}\cdot z_{S^{\prime}}\cdot a(S)\cdot a(S^{\prime})$
@@
The first inequlaity used that $\sum_{S^{\prime}\supseteq T}z_{S^{\prime}}\leq 1$ as $T\neq\emptyset$. Therefore, $M$ is PSD. ∎
Lemma 56 places a significant constraint on the distribution of triangles. We want to mention that the covariance matrix is implied by cluster LP, so we do not need to add extra constraints to the linear programming.
To leverage the property of the covariance matrix being positive semi-definite (PSD), we begin by partitioning the interval $[0,1]$ into numerous sub-intervals denoted as $I_{1},I_{2},...,I_{t}$ and each sub-interval $I_{j}\in[0,1]$ is continuous. Given any $y_{uv}\in[0,1]$, we use $I(y_{uv})$ to represent the interval containing $y_{uv}$. Let $l(I_{j})$ and $r(I_{j})$ be the left and right boundary of interval $I_{j}$. Given a node $u$, consider the matrix $Q_{u}\in\mathbb{R}^{t\times t}$, where $Q_{u}(I_{j},I_{k})$ is defined as
$Q_{u}(I_{j},I_{k})=\sum_{\begin{subarray}{c}y_{uv}\in I_{j},y_{uw}\in I_{k}\\ v\in V,w\in V\end{subarray}}y_{uvw}-y_{uv}y_{uw}$
@@
Specifically, a non-degenerate triangle $T$ increases the entries $Q_{u}(I(y_{uv}),I(y_{uw}))$ and $Q_{u}(I(y_{uw}),I(y_{uv}))$ by $y_{uvw}-y_{uv}y_{uw}$ for the matrix $Q_{u}$,. Similar computations are performed for $Q_{v}$ and $Q_{w}$. For degenerate triangle $T=(y_{uv},y_{uv},1,y_{uv})$, it increases $Q_{u}(I(y_{uv}),I(y_{uv}))$ by $(y_{uv}-y_{uv}^{2})$, increases $Q_{v}(I(y_{uv}),I(y_{uv}))$ by $(y_{uv}-y_{uv}^{2})$ and doesn’t affect and $Q_{w}$. In either case, we increase $Q(I(y_{uv}),I(y_{uw}))$ and $Q_{u}(I(y_{uw}),I(y_{uv}))$ by $y_{uvw}-y_{uv}y_{uw}$, $Q(I(y_{vu}),I(y_{vw}))$ and $Q(I(y_{vw}),I(y_{vu}))$ by $y_{uvw}-y_{vu}y_{vw}$, $Q(I(y_{wu}),I(y_{wv}))$ and $Q(I(y_{wv}),I(y_{wu}))$ by $y_{uvw}-y_{wu}y_{wv}$.
@@
$C^{(1)}(T)=C^{(2)}(T)=COV_{u}(v,w)=y_{uvw}-y_{uv}y_{uw}$
$C^{(3)}(T)=C^{(4)}(T)=COV_{v}(u,w)=y_{uvw}-y_{vu}y_{vw}$
$C^{(5)}(T)=C^{(6)}(T)=COV_{w}(v,u)=y_{uvw}-y_{wv}y_{wu}$
Let $Head(T)\in\mathbb{R}^{6\times t}$ be a binary matrix, defined as $Head(T)(i,I(a))=1$ when:
$(i,a)\in\{(1,y_{uv}),(2,y_{uw}),(3,y_{vu}),(4,y_{vw}),(5,y_{wu}),(6,y_{wv})\}$
@@
and 0 otherwise. Similarly, we define $Tail(T)\in\mathbb{R}^{6\times t}$ as a binary matrix, defined as $Tail(T)(i,I(a))=1$ when:
$(i,a)\in\{(1,y_{uw}),(2,y_{uv}),(3,y_{vw}),(4,y_{vu}),(5,y_{wv}),(6,y_{wu})\}$
We can then express $Q(y)$ as:
$Q(y)=\sum_{T\in\mathcal{T}}\eta_{T}(y)\cdot\left(Head^{\intercal}(T)\cdot C(T)\cdot Tail(T)\right)$
@@
Let $\mathcal{T}(I_{i},I_{j},I_{k})=\{(y_{uv},y_{uw},y_{vw},y_{uvw})\in\mathcal{T}\mid y_{uv}\leq y_{uw}\leq y_{vw},y_{uv}\in I_{i},y_{vw}\in I_{j},y_{uv}\in I_{k}\}$ be the set of triangles within the intervals $I_{i},I_{j},I_{k}$. For any two triangles $T$ with $(y_{uv}=a,y_{uw}=b,y_{vw}=c)$ and $T^{\prime}$ with $(y_{uv}=a^{\prime},y_{uw}=b^{\prime},y_{vw}=c^{\prime})$, if $T$ and $T^{\prime}$ are located within the same interval, meaning $I(a)=I(a^{\prime})$, $I(b)=I(b^{\prime})$, and $I(c)=I(c^{\prime})$, then we have $Head(T)=Head(T^{\prime})$ and $Tail(T)=Tail(T^{\prime})$. Given $I_{i},I_{j},I_{k}$, define $Head(I_{i},I_{j},I_{k})=Head(T)$ (or $Tail(I_{i},I_{j},I_{k})=Tail(T)$), where $T$ is an arbitrary triangle such that $T\in\mathcal{T}(I_{i},I_{j},I_{k})$. We can then express $Q(y)$ as:
$Q(y)=\sum_{I_{i},I_{j},I_{k}}Head^{\intercal}(I_{i},I_{j},I_{k})\cdot\left(\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot C(T)\right)\cdot Tail(I_{i},I_{j},I_{k})$
@@
Since $Q(y)\succeq 0$, it provides a constraint for $\eta_{T}(y)$. Our target is to show that $\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))\geq 0$ under the constraint that $Q(y)\succeq 0$. We will consider all possible triangles in each range $\mathcal{T}(I_{i},I_{j},I_{k})$ and there will be at most $O(t^{3})$ different range. For any solution $y$, the triangle $T=\{(y_{uv},y_{uw},y_{vw},y_{uvw})\mid y_{uv}\leq y_{uw}\leq y_{vw}\}$ satisfies the triangle inequality, that is $1-y_{uw}+1-y_{vw}\geq 1-y_{uv}$. Therefore, given range $I_{i},I_{j},I_{k}$, we assume that $l(I_{i})\geq l(I_{j})+l(I_{k})-1$, $r(I_{j})\leq 1+l(I_{i})-l(I_{k})$ and $r(I_{k})\leq 1+l(I_{i})-l(I_{j})$, since there won’t be any triangles in the range when the above inequality doesn’t hold. If necessary, we can truncate intervals to ensure that these inequalities hold true.
@@
In each range $I_{i},I_{j},I_{k}$, we will use $\tilde{d}(I_{i},I_{j},I_{k})=min_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}(\Delta(T)-\mathrm{cost}(T))$ to establish a lower bound on $\Delta(T)-\mathrm{cost}(T)$. $\tilde{d}(I_{i},I_{j},I_{k})$ can be precomputed and there will be at most $O(t^{3})$ different $\tilde{d}$ value. We first show that $\tilde{d}(I_{i},I_{j},I_{k})$ is a lower bound even for degenerate triangle. Recall that we treat degenerate triangle as triangle with $y$ value $(y_{uv},y_{uv},1,y_{uv})$.
@@
$cost(u,v)-\Delta(u,v)\geq\tilde{d}(I(y_{uv}),I(y_{uv}),I(1))$
@@
Let $T_{1}$ represent the ++++\!+\!++ + + triangle with $(y_{uv},y_{uv},1,y_{uv})$.If $(u,v)$ is a positive edge and $y_{uv}\geq 2/3$, then $cost(u,v)=0$ and $\Delta(u,v)=2b^{+}(x_{uv})$. In this case, $cost(T_{1})=0$ and $\Delta(T_{1})=2b^{+}(x_{uv})+b^{+}(0)$. If $(u,v)$ is a positive edge and $y_{uv}<2/3$, we have $cost(u,v)=2(1-y_{uv})$ and $\Delta(u,v)=2b^{+}(x_{uv})$. For $T_{1}$ in this scenario, $cost(T_{1})=2(1-y_{uv})$ and $\Delta(T_{1})=2b^{+}(x_{uv})+b^{+}(0)y_{uv}$. In either case, we observe that $\Delta(u,v)-cost(u,v)=\Delta(T_{1})-cost(T_{1})\geq\tilde{d}(I(y_{uv}),I(y_{uv}),I(1))$.
Let $T_{2}$ be the +-\!-\!+- - + triangle with $(y_{uv},y_{uv},1,y_{uv})$. If $(u,v)$ is a negative edge, then $cost(u,v)=2y_{uv}$ and $\Delta(u,v)=2b^{-}(x_{uv})$. For $T_{2}$, we have $cost(T_{2})=2y_{uv}+2y_{uv}-2y_{uv}^{2}$ and $\Delta(T_{2})=2b^{-}(x_{uv})+b^{+}(0)(2y_{uv}-y_{uv}^{2})$. Therefore, we conclude that $\Delta(u,v)-cost(u,v)\geq\Delta(T_{2})-cost(T_{2})\geq\tilde{d}(I(y_{uv}),I(y_{uv}),I(1))$.
We still need to represent $Q$ using $\eta_{T}$. The problem is there might be infinite types of triangles in each range $\mathcal{T}(I_{i},I_{j},I_{k})$, and we are not able to list every possible triangle. To solve this problem, our last observation is that, for $i\in[1,6]$, $C^{(i)}(T)$ is a multi-linear function in terms of $y_{uv},y_{uw},y_{vw}$ and $y_{uvw}$.
Since $C^{(i)}(T)$ is multi-linear, we only need to consider 16 different triangles in each range to cover all possible triangles, where those triangles’ $y$ value are
@@
$y_{uvw}$$\in\{max(0,(l(I_{i})+l(I_{j})+l(I_{k})-1)/2),r(I_{k})\}.$
The $y_{uvw}$ constraints comes from the fact that $y_{uvw}\leq y_{vw}$ and $y_{uw}+y_{uv}+y_{vw}-2y_{uvw}\leq 1$. Each triangle might be considered in different ranges multiple times and we shall treat them as different types of triangles. Let $\mathcal{T_{D}}(I_{i},I_{j},I_{k})$ be the set of triangles containing this 16 triangles in $\mathcal{T}(I_{i},I_{j},I_{k})$ and $\mathcal{T_{D}}=\cup_{I_{i},I_{j},I_{k}}\mathcal{T_{D}}(I_{i},I_{j},I_{k})$, we can set up a semi-definite programming to lower bound $\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))$. In factor-revealing SDP, the variables are $\eta_{T}$ that represent the ratio of triangles $T\in\mathcal{T_{D}}$ and $N\cdot\eta_{T}$ is the number of triangle of $T$, where $N=\frac{n(n-1)(n-2)}{6}+\frac{n(n-1)}{2}=\frac{n(n-1)(n+1)}{6}$. $\tilde{d}(I_{i},I_{j},I_{k})$, $Head^{\intercal}(I_{i},I_{j},I_{k})$, $C(T)$ and $Tail(I_{i},I_{j},I_{k})$ are constant and can be pre-computed.
[…]
@@
Note that $\tilde{d}(I_{i},I_{j},I_{k})$ is the lower bound of $\Delta(T)-\mathrm{cost}(T)$ for any $T\in\mathcal{T}(I_{i},I_{j},I_{k})$. Hence, we have:
The difficulty of the proof arises from the fact that we only consider triangles from $\mathcal{T_{D}}(I_{i},I_{j},I_{k})$, which contains only 16 triangles. We will show that we can use triangles from $\mathcal{T_{D}}$ to represent $Q$ and the objective value is at most $\sum_{I_{i},I_{j},I_{k}}\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot\tilde{d}(I_{i},I_{j},I_{k})$. For any triangle $T$ with $(y_{uv}\in I_{i},y_{uw}\in I_{j},y_{vw}\in I_{k})$ and $l(I_{i})\leq l(I_{j})\leq l(I_{k})$, let $T_{1},T_{2},...,T_{16}$ be the 16 triangles from $\mathcal{T_{D}}(I_{i},I_{j},I_{k})$. We will later show that there exist $\lambda_{i}\geq 0$ such that $C(T)=\sum_{i\in[1,16]}\lambda_{i}\cdot C(T_{i})$ and $\sum_{i\in[1,16]}\lambda_{i}=1$. Therefore, we can use $T_{1},T_{2},...,T_{16}$ to replace $T$ without affecting the $Q$ value. As we use $\tilde{d}(I_{i},I_{j},I_{k})$ in the objective function and $\sum_{i\in[1,16]}\lambda_{i}=1$, this substitution doesn’t impact the objective value either.
@@
$\geq\sum_{I_{i},I_{j},I_{k}}\sum_{T\in\mathcal{T_{D}}(I_{i},I_{j},I_{k})}N\cdot\eta_{T}\cdot\tilde{d}(I_{i},I_{j},I_{k})\geq N\cdot{\mathrm{OPT}_{\mathrm{SDP}}}$
@@
Theorem 59 provides a valuable method for establishing a lower bound on $\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))$ for any given solution and budget function. To proceed with the computation of this lower bound, we need to solve factor-revealing SDP. It’s important to emphasize that solving factor-revealing SDP is completely independent of cluster LP. The value of ${\mathrm{OPT}_{\mathrm{SDP}}}$ is determined solely by the chosen budget function and the selected intervals $I_{1},I_{2},...,I_{t}$.
@@
$Head^{\intercal}(T)\cdot A\cdot Tail(T)$
It’s worth noting this increase holds even for degenerate triangles. For any degenerate triangle $(u,v)$, it raises $F(I(y_{uv}),I(y_{uv}))$, $F(I(y_{uv}),I(1))$, and $F(I(1),I(y_{uv}))$ by 2, echoing the increase seen in the triangle with the $y$ value $(y_{uv},y_{uv},1,y_{uv})$. However, unlike the $Q_{u}$ vector, $F_{u}$ also considers cases where $v=u$ and $w=u$. In the $Q_{u}$ vector, this case doesn’t impact $Q_{u}$ due to $y_{uuu}-y_{uu}y_{uu}=0$. We need to increment $F(I(1),I(1))$ by $n$ since $\eta(y)$ does not account for $v=u$ and $w=u$. Fortunately, we already include the ++++\!+\!++ + + triangle $T_{3}=(1,1,1,1)$; given that $T_{3}$ doesn’t influence the matrix $Q$ value and $\Delta(T_{3})-cost(T_{3})=0$, each occurrence of $T_{3}$ raises $F(I(1),I(1))$ by 6. By increasing $\eta_{T_{3}}(y)$ by $n/6$, we can effectively represent $F$ as:
$F(y)=\sum_{I_{i},I_{j},I_{k}}Head^{\intercal}(I_{i},I_{j},I_{k})\left(\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot A\right)\cdot Tail(I_{i},I_{j},I_{k})$

### v3

@@ 4.3 Using Independently Sampled Clusters to Construct Cluster LP Solution
Given cluster LP solution $z_{S}$ and $y$, for each node $u$, the covariance matrix $\mathrm{COV}_{u}\in\mathbb{R}^{V\times V}$ where
$\mathrm{COV}_{u}(v,w)=y_{uvw}-y_{uv}y_{uw}$
is positive semidefinite (PSD). Actually, we can even show a stronger version of the covariance constraint.
@@
$\geq\sum_{S\supseteq T,S^{\prime}\supseteq T}z_{S}z_{S^{\prime}}\cdot a^{2}(S)-\sum_{S\supseteq T}\sum_{S^{\prime}\supseteq T}z_{S}\cdot z_{S^{\prime}}\cdot a(S)\cdot a(S^{\prime})$
@@
The first inequality used that $\sum_{S^{\prime}\supseteq T}z_{S^{\prime}}\leq 1$ as $T\neq\emptyset$. Therefore, $M$ is PSD. ∎
Lemma 48 places a significant constraint on the distribution of triangles. We want to mention that the covariance matrix is implied by cluster LP, so we do not need to add extra constraints to the linear program. To obtain Lemma 48, we only require that the solution is from cluster LP.
To leverage the property of the covariance matrix being positive semidefinite, we begin by partitioning the interval $[0,1]$ into numerous subintervals denoted as $I_{1},I_{2},...,I_{t}$ where each subinterval $I_{j}\in[0,1]$ is continuous. Given any $y_{uv}\in[0,1]$, we use $I(y_{uv})$ to represent the interval containing $y_{uv}$. Let $l(I_{j})$ and $r(I_{j})$ be the left and right boundary of interval $I_{j}$. Given a node $u$, consider the matrix $Q_{u}\in\mathbb{R}^{t\times t}$, where $Q_{u}(I_{j},I_{k})$ is defined as
$Q_{u}(I_{j},I_{k})=\sum_{\begin{subarray}{c}y_{uv}\in I_{j},y_{uw}\in I_{k}\\ v\in V,w\in V\end{subarray}}y_{uvw}-y_{uv}y_{uw}.$
@@
Specifically, a non-degenerate triangle $T$ increases the entries $Q_{u}(I(y_{uv}),I(y_{uw}))$ and $Q_{u}(I(y_{uw}),I(y_{uv}))$ by $y_{uvw}-y_{uv}y_{uw}$ for the matrix $Q_{u}$,. Similar computations are performed for $Q_{v}$ and $Q_{w}$. For degenerate triangle $T=(y_{uv},y_{uv},1,y_{uv})$, it increases $Q_{u}(I(y_{uv}),I(y_{uv}))$ by $(y_{uv}-y_{uv}^{2})$, increases $Q_{v}(I(y_{uv}),I(y_{uv}))$ by $(y_{uv}-y_{uv}^{2})$ and does not affect $Q_{w}$. In either case, we increase $Q(I(y_{uv}),I(y_{uw}))$ and $Q_{u}(I(y_{uw}),I(y_{uv}))$ by $y_{uvw}-y_{uv}y_{uw}$, $Q(I(y_{vu}),I(y_{vw}))$ and $Q(I(y_{vw}),I(y_{vu}))$ by $y_{uvw}-y_{vu}y_{vw}$, $Q(I(y_{wu}),I(y_{wv}))$ and $Q(I(y_{wv}),I(y_{wu}))$ by $y_{uvw}-y_{wu}y_{wv}$.
@@
$C^{(1)}(T)=C^{(2)}(T)=\mathrm{COV}_{u}(v,w)=y_{uvw}-y_{uv}y_{uw},$
$C^{(3)}(T)=C^{(4)}(T)=\mathrm{COV}_{v}(u,w)=y_{uvw}-y_{vu}y_{vw},$
$C^{(5)}(T)=C^{(6)}(T)=\mathrm{COV}_{w}(v,u)=y_{uvw}-y_{wv}y_{wu}.$
Let $Head(T)\in\mathbb{R}^{6\times t}$ be a binary matrix, defined as $Head(T)(i,I(a))=1$ when
$(i,a)\in\{(1,y_{uv}),(2,y_{uw}),(3,y_{vu}),(4,y_{vw}),(5,y_{wu}),(6,y_{wv})\},$
@@
and 0 otherwise. Similarly, we define $Tail(T)\in\mathbb{R}^{6\times t}$ as a binary matrix, defined as $Tail(T)(i,I(a))=1$ when
$(i,a)\in\{(1,y_{uw}),(2,y_{uv}),(3,y_{vw}),(4,y_{vu}),(5,y_{wv}),(6,y_{wu})\}.$
We can then express $Q(y)$ as
$Q(y)=\sum_{T\in\mathcal{T}}\eta_{T}(y)\cdot\left(Head^{\intercal}(T)\cdot C(T)\cdot Tail(T)\right).$
@@
Let $\mathcal{T}(I_{i},I_{j},I_{k})=\{(y_{uv},y_{uw},y_{vw},y_{uvw})\in\mathcal{T}\mid y_{uv}\leq y_{uw}\leq y_{vw},y_{uv}\in I_{i},y_{vw}\in I_{j},y_{uv}\in I_{k}\}$ be the set of triangles within the intervals $I_{i},I_{j},I_{k}$. For any two triangles $T$ with $(y_{uv}=a,y_{uw}=b,y_{vw}=c)$ and $T^{\prime}$ with $(y_{uv}=a^{\prime},y_{uw}=b^{\prime},y_{vw}=c^{\prime})$, if $T$ and $T^{\prime}$ are located within the same interval, meaning $I(a)=I(a^{\prime})$, $I(b)=I(b^{\prime})$, and $I(c)=I(c^{\prime})$, then we have $Head(T)=Head(T^{\prime})$ and $Tail(T)=Tail(T^{\prime})$. Given $I_{i},I_{j},I_{k}$, define $Head(I_{i},I_{j},I_{k})=Head(T)$ (or $Tail(I_{i},I_{j},I_{k})=Tail(T)$), where $T$ is an arbitrary triangle such that $T\in\mathcal{T}(I_{i},I_{j},I_{k})$. We can then express $Q(y)$ as
$Q(y)=\sum_{I_{i},I_{j},I_{k}}Head^{\intercal}(I_{i},I_{j},I_{k})\cdot\left(\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot C(T)\right)\cdot Tail(I_{i},I_{j},I_{k}).$
@@
Since $Q(y)\succeq 0$, it provides a constraint for $\eta_{T}(y)$. Our target is to show that $\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))\geq 0$ under the constraint that $Q(y)\succeq 0$. We will consider all possible triangles in each range $\mathcal{T}(I_{i},I_{j},I_{k})$ and there will be at most $O(t^{3})$ different ranges. For any solution $y$, the triangle $T=\{(y_{uv},y_{uw},y_{vw},y_{uvw})\mid y_{uv}\leq y_{uw}\leq y_{vw}\}$ satisfies the triangle inequality, that is $1-y_{uw}+1-y_{vw}\geq 1-y_{uv}$. Therefore, given range $I_{i},I_{j},I_{k}$, we assume that $l(I_{i})\geq l(I_{j})+l(I_{k})-1$, $r(I_{j})\leq 1+l(I_{i})-l(I_{k})$ and $r(I_{k})\leq 1+l(I_{i})-l(I_{j})$, since there will not be any triangles in the range when the above inequality does not hold. If necessary, we can truncate intervals to ensure that these inequalities hold true.
@@
In each range $I_{i},I_{j},I_{k}$, we will use $\tilde{d}(I_{i},I_{j},I_{k})=\min_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}(\Delta(T)-\mathrm{cost}(T))$ to establish a lower bound on $\Delta(T)-\mathrm{cost}(T)$. $\tilde{d}(I_{i},I_{j},I_{k})$ can be precomputed and there will be at most $O(t^{3})$ different $\tilde{d}$ value. We first show that $\tilde{d}(I_{i},I_{j},I_{k})$ is a lower bound even for degenerate triangle. Recall that we treat degenerate triangle as triangle with $y$ value $(y_{uv},y_{uv},1,y_{uv})$.
@@
$\mathrm{cost}(u,v)-\Delta(u,v)\geq\tilde{d}(I(y_{uv}),I(y_{uv}),I(1))$
@@
Let $T_{1}$ represent the ++++\!+\!+ triangle with $(y_{uv},y_{uv},1,y_{uv})$. If $(u,v)$ is a positive edge and $y_{uv}\geq 2/3$, then $\mathrm{cost}(u,v)=0$ and $\Delta(u,v)=2\cdot\mathrm{budget}^{+}(x_{uv})$. In this case, $\mathrm{cost}(T_{1})=0$ and $\Delta(T_{1})=2\cdot\mathrm{budget}^{+}(x_{uv})+\mathrm{budget}^{+}(0)$. If $(u,v)$ is a positive edge and $y_{uv}<2/3$, we have $\mathrm{cost}(u,v)=2(1-y_{uv})$ and $\Delta(u,v)=2\cdot\mathrm{budget}^{+}(x_{uv})$. For $T_{1}$ in this scenario, $\mathrm{cost}(T_{1})=2(1-y_{uv})$ and $\Delta(T_{1})=2\cdot\mathrm{budget}^{+}(x_{uv})+\mathrm{budget}^{+}(0)y_{uv}$. In either case, we observe that $\Delta(u,v)-\mathrm{cost}(u,v)=\Delta(T_{1})-\mathrm{cost}(T_{1})\geq\tilde{d}(I(y_{uv}),I(y_{uv}),I(1))$.
Let $T_{2}$ be the +-\!-\!+ triangle with $(y_{uv},y_{uv},1,y_{uv})$. If $(u,v)$ is a negative edge, then $\mathrm{cost}(u,v)=2y_{uv}$ and $\Delta(u,v)=2\cdot\mathrm{budget}^{-}(x_{uv})$. For $T_{2}$, we have $\mathrm{cost}(T_{2})=2y_{uv}+2y_{uv}-2y_{uv}^{2}$ and $\Delta(T_{2})=2\cdot\mathrm{budget}^{-}(x_{uv})+b^{+}(0)(2y_{uv}-y_{uv}^{2})$. Therefore, we conclude that $\Delta(u,v)-\mathrm{cost}(u,v)\geq\Delta(T_{2})-\mathrm{cost}(T_{2})\geq\tilde{d}(I(y_{uv}),I(y_{uv}),I(1))$. ∎
We still need to represent $Q$ using $\eta_{T}$. The problem is there might be infinite types of triangles in each range $\mathcal{T}(I_{i},I_{j},I_{k})$, and we are not able to list every possible triangle. To solve this problem, our last observation is that, for $i\in[1,6]$, $C^{(i)}(T)$ is a multilinear function in terms of $y_{uv},y_{uw},y_{vw}$ and $y_{uvw}$.
Since $C^{(i)}(T)$ is multilinear, we only need to consider 16 different triangles in each range to cover all possible triangles, where those triangles’ $y$ value are
@@
$y_{uvw}\in\{\max\Bigl(0,l(I_{i})+l(I_{j})-1$ $,l(I_{j})+l(I_{k})-1,l(I_{i})+l(I_{k})-1\Bigl),\min\Bigl(r(I_{i}),r(I_{j}),r(I_{k})\Bigl)\}.$
The $y_{uvw}$ constraints comes from the fact that $y_{uvw}\leq y_{vw}$ and $y_{uw}+y_{uv}-y_{uvw}\leq 1$. Each triangle might be considered in different ranges multiple times and we shall treat them as different types of triangles. Let $\mathcal{T_{D}}(I_{i},I_{j},I_{k})$ be the set of triangles containing this 16 triangles in $\mathcal{T}(I_{i},I_{j},I_{k})$ and $\mathcal{T_{D}}=\cup_{I_{i},I_{j},I_{k}}\mathcal{T_{D}}(I_{i},I_{j},I_{k})$, we can set up a semidefinite programming to lower bound $\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))$. In factor-revealing SDP, the variables are $\eta_{T}$ that represent the ratio of triangles $T\in\mathcal{T_{D}}$ and $N\cdot\eta_{T}$ is the number of triangle of $T$, where $N=\frac{n(n-1)(n-2)}{6}+\frac{n(n-1)}{2}=\frac{n(n-1)(n+1)}{6}$. $\tilde{d}(I_{i},I_{j},I_{k})$, $Head^{\intercal}(I_{i},I_{j},I_{k})$, $C(T)$ and $Tail(I_{i},I_{j},I_{k})$ are constant and can be precomputed.
$\min\qquad\sum_{I_{i},I_{j},I_{k}}\sum_{T\in\mathcal{T_{D}}(I_{i},I_{j},I_{k})}\eta_{T}\cdot\tilde{d}(I_{i},I_{j},I_{k})\qquad\text{s.t.}$ (factor-revealing SDP)
@@
Note that $\tilde{d}(I_{i},I_{j},I_{k})$ is the lower bound of $\Delta(T)-\mathrm{cost}(T)$ for any $T\in\mathcal{T}(I_{i},I_{j},I_{k})$. Hence, we have
$\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))\geq\sum_{I_{i},I_{j},I_{k}}\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot\tilde{d}(I_{i},I_{j},I_{k}).$
The difficulty of the proof arises from the fact that we only consider triangles from $\mathcal{T_{D}}(I_{i},I_{j},I_{k})$, which contains only 16 triangles. We will show that we can use triangles from $\mathcal{T_{D}}$ to represent $Q$ and the objective value is at most $\sum_{I_{i},I_{j},I_{k}}\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot\tilde{d}(I_{i},I_{j},I_{k})$. For any triangle $T$ with $(y_{uv}\in I_{i},y_{uw}\in I_{j},y_{vw}\in I_{k})$ and $l(I_{i})\leq l(I_{j})\leq l(I_{k})$, let $T_{1},T_{2},...,T_{16}$ be the 16 triangles from $\mathcal{T_{D}}(I_{i},I_{j},I_{k})$. We will later show that there exist $\lambda_{i}\geq 0$ such that $C(T)=\sum_{i\in[1,16]}\lambda_{i}\cdot C(T_{i})$ and $\sum_{i\in[1,16]}\lambda_{i}=1$. Therefore, we can use $T_{1},T_{2},...,T_{16}$ to replace $T$ without affecting the $Q$ value. As we use $\tilde{d}(I_{i},I_{j},I_{k})$ in the objective function and $\sum_{i\in[1,16]}\lambda_{i}=1$, this substitution does not impact the objective value either.
@@
$\geq\sum_{I_{i},I_{j},I_{k}}\sum_{T\in\mathcal{T_{D}}(I_{i},I_{j},I_{k})}N\cdot\eta_{T}\cdot\tilde{d}(I_{i},I_{j},I_{k})\geq N\cdot{\mathrm{OPT}_{\mathrm{SDP}}}.$
@@
Theorem 51 provides a valuable method for establishing a lower bound on $\sum_{T}\eta_{T}(y)\cdot(\Delta(T)-\mathrm{cost}(T))$ for any given solution and budget function. To proceed with the computation of this lower bound, we need to solve factor-revealing SDP. It is important to emphasize that solving factor-revealing SDP is completely independent of cluster LP. The value of $\mathrm{OPT}_{\mathrm{SDP}}$} is determined solely by the chosen budget function and the selected intervals $I_{1},I_{2},...,I_{t}$.
@@
$Head^{\intercal}(T)\cdot A\cdot Tail(T).$
It is worth noting this increase holds even for degenerate triangles. For any degenerate triangle $(u,v)$, it raises $F(I(y_{uv}),I(y_{uv}))$, $F(I(y_{uv}),I(1))$, and $F(I(1),I(y_{uv}))$ by 2, echoing the increase seen in the triangle with the $y$ value $(y_{uv},y_{uv},1,y_{uv})$. However, unlike the $Q_{u}$ vector, $F_{u}$ also considers cases where $v=u$ and $w=u$. In the $Q_{u}$ vector, this case does not impact $Q_{u}$ due to $y_{uuu}-y_{uu}y_{uu}=0$. We need to increment $F(I(1),I(1))$ by $n$ since $\eta(y)$ does not account for $v=u$ and $w=u$. Fortunately, we already include the ++++\!+\!+ triangle $T_{3}=(1,1,1,1)$; given that $T_{3}$ does not influence the matrix $Q$ value and $\Delta(T_{3})-\mathrm{cost}(T_{3})=0$, each occurrence of $T_{3}$ raises $F(I(1),I(1))$ by 6. By increasing $\eta_{T_{3}}(y)$ by $n/6$, we can effectively represent $F$ as:
$F(y)=\sum_{I_{i},I_{j},I_{k}}Head^{\intercal}(I_{i},I_{j},I_{k})\left(\sum_{T\in\mathcal{T}(I_{i},I_{j},I_{k})}\eta_{T}(y)\cdot A\right)\cdot Tail(I_{i},I_{j},I_{k}).$